Cbenavidez:

The equilibrium constant for each reaction of lead sulfate dissolving in water is 1.5 x 10-8. If enough PbS04 is added for equilibrium to be reached what will be the equilibrium concentration of Pb2=?

WickedHeart:

Hi, what would you say the answer would be? @cbenavidez

Cbenavidez:

Hi I picked C. [pb^2+][so4^2-] 1.5 X10^-8 1.5X10^-8=x^2 I squared both sides but something doesn't seem right. My options are: A. 2.25 x 10-16M B. 7.0 x 10-9M C. 1.22 x 10-4M D. This cannot be determined without knowing the initial concentration of PBS04.

WickedHeart:

Keq = [Pb^2+][SO4^2-] PbSO4 is not important as it is a solid and it will not affect the equation we know Keq= 1.5x10^-8 1.5x10^-8 = [x][x] 1.5x10^-8 = x^2 square both sides and use the positive x value for your answer. x is equivalent to both Pb^2+ and SO4^2+ so you have found the concentration of both with one calculation.

WickedHeart:

what would you get?

Cbenavidez:

Honestly, I don't know. I keep coming up with different answers. I even came up with 2.25x 10-16M

Rosee5656:

.